# Revisiting 360 Depth Estimation with PanoGabor: A New Fusion Perspective

### v1

@@ IV-A Evaluation Metrics
TABLE II: Ablation study.
@@
IV-B Comparison Analysis
Quantitative Results. We exhibit the quantitative comparison results in Table I, which demonstrate that our model outperforms all other state-of-the-art solutions on most metrics. Our PGFuse method significantly outperforms compared methods. Specifically, on the real-world Matterport3D and Stanford2D3D datasets, our approach demonstrates notable improvements in the important RMSE and $\delta_{1}$ metrics. Additionally, it achieves comparable results on the virtual Structured3D dataset, showcasing its robustness and effectiveness across various environments. These results highlight the efficacy of PGFuse in monocular 360-depth estimation across diverse and challenging datasets. Although our performance on the loosest accuracy metric $\delta_{3}$ is slightly lower than the Elite360D [7] and the Unifuse [2], it performs better than the two methods on the other two tighter accuracy metrics ($\delta_{1}$ and $\delta_{2}$). Note that in the small-scale realistic Standford2D3D dataset, PGFuse outperforms other methods in all metrics. In [7], the authors argue that due to the limited data in the Stanford2D3D data set, the transformer-based approach can hardly provide accurate global perception, while the core components of our method are CNN-based and cannot be influenced by this negative effect. As for the Structred3D dataset, we can observe that our approach performs badly on the Sq Rel metric than the Unifuse and Elite360D, we suspect it might be related to the pre-processing operation. For a fair comparison, we follow the Elite360D to process the data (normalize the data and scale it to 10m), which confuses the model in estimating an accurate depth range.
[…]
Complexity. We show the complexity comparison results in Table I. We can observe that those transformer-based architectures may have fewer parameters and FLOPs. However, the complexity of these approaches is quadric to the input pixels, which limits its application to high-resolution images. In addition, their graphics memory cost is huge, which makes it difficult to use in practice. The other SoTA approach, Bifuse, has the largest parameters and huge FLOPs. And the training time is unbearable for large-scale datasets. Although our scheme has slightly more parameters and FLOPs than the Elite360D, the proposed approach performs better than Elite360D, especially on the RMSE metric, by a significant margin. In addition, we argue that the core fusion component of the Elite360D is attention-based, similar to the transformer-based frameworks, resulting in fewer parameters and FLOPs.
Figure 6: Visualization of the heatmap of the features before/after the CS-UFM module. Because each face in the cubemap projection is processed individually, resulting in discontinuous ERP features. The CS-UFM module can address this issue effectively.
@@ IV-C Ablation Study
To demonstrate the effectiveness of the main proposed components, we follow Unifuse [2] and Elite360D [7] to design the ablation study on the Matterport3D dataset [9]. All the models are with the same conditions and trained to converge. As illustrated in Figure 2, we remove the fusion stage from the full framework to serve as the baseline for conducting the ablation study. The experimental results of the baseline are presented as ‘a’ in Table II. We further add the proposed CS-UFM module, a version without the PanoGabor layer embedded as ‘b’. Index ‘c’ refers to the use of the full CS-UFM module. Index ‘d’ indicates that we employ the proposed spherical gradient constraint to stabilize the orientation sensitivity of the Gabor filters. Additionally, in ‘e’, all PanoGabor layers are replaced with normal Gabor filters to further verify the effectiveness of the designed PanoGarbor.
@@
Effectiveness of the CS-UFM module. From Table II, we can observe that the CS-UFM module can obtain an improvement even without the designed PanoGabor layer. This is because the fusion module can effectively fuse the two sets of features due to the channel-wise and spatial-wise designs. Moreover, when further embedding the proposed PanoGabor layer, our approach gets a significant improvement across all the evaluation metrics, verifying the necessity of addressing the reintroduced distortion problems during the fusion process.
@@
Effectiveness of the PanoGabor layer. Comparing the b and e in Table II, we can observe that benefiting from the frequency domain transformation along with the capability of reducing surface texture interference, the normal Gabor layer can benefit the fusion process. From d and e, we can observe that our designed PanoGabor performs better than the normal ones due to the distortion-aware designs.
@@
Effectiveness of the spherical gradient constraint. Recently, some works [4, 7] prefer to employ a gradient-based constraint to sharpen the depth edges. This strategy aims to extract the horizontal/vertical structures to design the objective function. However, since the Gabor filters are orientation-sensitive, we design the spherical gradient constraint to incorporate the panoramic geometry to stabilize this orientation sensitivity. From Table II, we can observe that the PanoGabor with the proposed spherical gradient constraint performs much better than that without.
@@
Various Backbones. Following Elite360D [7], we also adopt various backbones for our framework. From Table III, we can observe that our scheme can benefit from employing more powerful backbones.
@@
Trade-off between performance and efficiency. From Table IV, Elite360D has obtained a significant improvement by employing EfficientNet-B5 as a feature extractor. However, it brings unbearable memory costs. In contrast, our approach can get a superior performance just by replacing a group of input features without additional memory cost. This also demonstrates the robustness and scalability of our PGFuse.
TABLE III: Ablation study on various backbones.
@@
Datasets Backbones Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$
ResNet-18 0.1125 0.0977 0.4933 87.78
M3D ResNet-34 0.1067 0.0858 0.4624 89.06
ResNet-50 0.1058 0.0841 0.4618 88.93
@@
TABLE IV: More advanced performance comparison results. * represent that Elite360D employs Efficient-B5 as backbone. indicates that we replace one of the cubemap features with extracted ERP features.
@@
Datasets Method Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$
M3D Elite360D* 0.1048 0.0805 0.4524 89.92
M3D Ours 0.0972 0.0801 0.4480 90.67
@@ V Conclusion
In this paper, we propose a Gabor-based framework, termed PGFuse for 360 depth estimation. Since the existing fuse-based scheme ignores the reintroduced distortion when integrating other representations into ERP format, our proposed PGFuse embeds a Gabor-based layer to a unidirectional fusion module to explicitly deal with the reintroduced distortion. Specifically, to make the Gabor filters adapt to 360 images, we first introduce a linear latitude-aware representation to model the panoramic distortion distribution pattern. Then we leverage this representation to generate customized Gabor filters. These tuned PanoGabor can enhance the depth cues in the frequency domain and achieve effective distortion awareness. Therefore, our fusion module can deliver spatial-continuous and content-consistent ERP features. Extensive experiment results demonstrate that our approach outperforms current SoTA approaches.

### v4

@@ IV-A Evaluation Metrics
IV-B Comparison Analysis
Quantitative Results. We exhibit the quantitative comparison results in Table LABEL:tab:quanti, which demonstrate that our model outperforms all other state-of-the-art solutions on most metrics. Our PGFuse method significantly outperforms compared methods. Specifically, on the real-world Matterport3D and Stanford2D3D datasets, our approach demonstrates notable improvements in the important RMSE and $\delta_{1}$ metrics. Additionally, it achieves comparable results on the virtual Structured3D dataset, showcasing its robustness and effectiveness across various environments. These results highlight the efficacy of PGFuse in monocular 360-depth estimation across diverse and challenging datasets. Although our performance on the loosest accuracy metric $\delta_{3}$ is slightly lower than the Elite360D [7] and the Unifuse [2], it performs better than the two methods on the other two tighter accuracy metrics ($\delta_{1}$ and $\delta_{2}$). Note that in the small-scale realistic Standford2D3D dataset, PGFuse outperforms other methods in all metrics. In [7], the authors argue that due to the limited data in the Stanford2D3D dataset, the transformer-based approach can hardly provide accurate global perception, while the core components of our method are CNN-based and cannot be influenced by this negative effect. As for the Structred3D dataset, we can observe that our approach performs poorly on the Sq Rel metric than the Unifuse and Elite360D, we suspect it might be related to the pre-processing operation. For a fair comparison, we follow the Elite360D to process the data (normalize the data and scale it to 10m), which confuses the model in estimating an accurate depth range.
[…]
Complexity. We show the complexity comparison results in Table II. We can observe that those transformer-based architectures may have fewer parameters and FLOPs. However, the complexity of these approaches is quadric to the input pixels, which limits its application to high-resolution images. In addition, their graphics memory cost is huge, which makes it difficult to use in practice. The other SoTA approach, Bifuse, has the largest parameters and huge FLOPs. And the training time is unbearable for large-scale datasets. Although our scheme has slightly more parameters and FLOPs than the Elite360D, the proposed approach performs better than Elite360D, especially on the RMSE metric, by a significant margin. In addition, we argue that the core fusion component of the Elite360D is attention-based, similar to the transformer-based frameworks, resulting in fewer parameters and FLOPs.
TABLE III: Ablation study.
@@ IV-B Comparison Analysis
  • The bold values indicate the best performance.
Figure 9: Visualization of the heatmap of the features w/o and w/ the CS-UFM module. Because each face in the cubemap projection is processed individually, resulting in discontinuous ERP features. The CS-UFM module can eliminate this issue effectively.
Figure 10: Qualitative comparison results of ablation study. The areas marked in red rectangles indicate that our approach gains superior performance.
@@ IV-C Ablation Study
To demonstrate the effectiveness of the main proposed components, we follow Unifuse [2] and Elite360D [7] to design the ablation study on the Matterport3D dataset [13]. All the models have the same conditions and are trained to converge. As illustrated in Fig. 2, we remove the fusion stage from the full framework to serve as the baseline for conducting the ablation study. The experimental results of the baseline are presented as ‘a’ in Table LABEL:tab:ablation. We further add the proposed CS-UFM module, a version without the PanoGabor layer embedded as ‘b’. Index ‘c’ refers to the use of the full CS-UFM module. Index ‘d’ indicates that we employ the proposed spherical gradient constraint to stabilize the orientation sensitivity of the Gabor filters. Additionally, in ‘e’, all PanoGabor layers are replaced with normal Gabor filters to further verify the effectiveness of the designed PanoGarbor.
@@
Effectiveness of the CS-UFM module. From Table LABEL:tab:ablation and Fig. 10, we can observe that the CS-UFM module can obtain an improvement even without the designed PanoGabor layer. This is because the fusion module can effectively fuse the two sets of features due to the channel-wise and spatial-wise designs. Moreover, when further embedding the proposed PanoGabor layer, our approach gets a significant improvement across all the evaluation metrics, verifying the necessity of addressing the reintroduced distortion problems during the fusion process. From Fig. 9, we can observe that our CS-UFM module has a uniformly distributed heat map, indicating that it can effectively fuse features and handle the reintroduced distortion.
@@
Effectiveness of the PanoGabor layer. Comparing the b and e in Table LABEL:tab:ablation and Fig. 10, we can observe that benefiting from the frequency domain transformation along with the capability of reducing surface texture interference, the normal Gabor layer can benefit the fusion process. From d and e, we can observe that our designed PanoGabor performs better than the normal ones due to the distortion-aware designs.
@@
Effectiveness of the spherical gradient constraint. Recently, some works [4, 7] prefer to employ a gradient-based constraint to sharpen the depth edges. This strategy aims to extract the horizontal/vertical structures to design the objective function. However, since the Gabor filters are orientation-sensitive, we design the spherical gradient constraint to incorporate the panoramic geometry to stabilize this orientation sensitivity. From Table LABEL:tab:ablation and Fig. 10, we can observe that the PanoGabor with the proposed spherical gradient constraint performs much better than that without.
@@
Various Backbones. Following Elite360D [7], we also adopt various backbones for our framework. From Table IV, we can observe that our scheme can benefit from employing more powerful backbones.
@@
Trade-off between performance and efficiency. From Table V, Elite360D has obtained a significant improvement by employing EfficientNet-B5 as a feature extractor. However, it brings unbearable memory costs. In contrast, our approach can achieve superior performance just by replacing a group of input features without additional memory cost. This also demonstrates the robustness and scalability of our PGFuse.
TABLE IV: Ablation study on various backbones.
@@
[b] Datasets Backbone Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$ ResNet-18 0.1125 0.0977 0.4933 87.78 M3D ResNet-34 0.1067 0.0858 0.4624 89.06 ResNet-50 0.1058 0.0841 0.4618 88.93
  • The bold values indicate the best performance.
@@
TABLE V: More advanced performance comparison results. * represent that Elite360D employs Efficient-B5 as backbone. indicates that we replace one of the cubemap features with extracted ERP features.
@@
[b] Datasets Method Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\mathrm{RMSE}\downarrow$ $\delta_{1}(\%)\uparrow$ M3D Elite360D* 0.1048 0.0805 0.4524 89.92 M3D Ours 0.0972 0.0801 0.4480 90.67
  • The bold values indicate the best performance.
@@ V Conclusion
In this paper, we have proposed a Gabor-based framework, termed PGFuse for 360 depth estimation. Since the existing fuse-based scheme ignores the reintroduced distortion when integrating other representations into ERP format, our proposed PGFuse embeds a Gabor-based layer into a unidirectional fusion module to explicitly deal with the reintroduced distortion. Specifically, to make the Gabor filters adapt to 360 images, we first introduce a latitude-aware representation to model the panoramic distortion distribution pattern. Then we leverage this representation to generate customized Gabor filters. These tuned PanoGabor can enhance the depth cues (i.e., resistant to surface texture and lighting interference, while preserving structural details) in the frequency domain and achieve effective distortion awareness. Therefore, our fusion module can deliver spatial-continuous and content-consistent ERP features. Extensive experiment results demonstrate that our approach outperforms current SoTA approaches.